OpenStudy (anonymous):

The equation of line WX is y = −2x − 5. Write an equation of a line perpendicular to line WX in slope-intercept form that contains point (−1, −2).

OpenStudy (anonymous):

Y=mx+b M=slope

OpenStudy (anonymous):

for a perpendicular line the slope would be the negative reciprocal of the slope in the given equation so the slope of the perpendicular like would be m=1/2 so you would have y=1/2x+b

OpenStudy (anonymous):

Then you would have to plug in the points given into the x and y so solve for b so you would have -2=(1/2)(-1)+b

OpenStudy (anonymous):

the equation should be y=1/2x+the answer for b :)

OpenStudy (anonymous):

Thanks !!

OpenStudy (anonymous):

No problem!!